In a circle of radius 17, find the length of the arc spanned by a central angle pi/4 radians. Use 3.1416 for pi and round to 2 decimal places. Help please I am not understanding!!
Answer by Cromlix(4381)

Formula = Angle/360 = length of arc/ Pi*Diameter = area of sector/Pi*r^2
Using the first two
Angle/360 = length of arc/ Pi* D
Pi/4 = 45 degrees
45/360 = length of arc/Pi * 34
45 * Pi * 34/ 360 = length of arc
4806.648/360 = length of arc
13.3518 = length of arc
13.35 units.
Hope this helps.
:-)